OpenStudy (anonymous):

On a map’s coordinate grid, Panthersville is located at (-3, 2), and Heel City is located at (4, 8). Falconton is the midpoint between Panthersville and Heel City. What is the approximate distance from Panthersville to Falconton? (One map unit equals one mile.)

OpenStudy (anonymous):

Get out graph paper and actually set it up. What you can do is set up right triangles to show distances. For example, Panthersville on the x is at -3 and Heel is at 4. The distance between the x's is 7. The distance between their y's is 6. If you use pythagorean's theorem, (a^2 + b^2 = c^2), and put in 49 +36 = c^2, the distance is root 85 i think

OpenStudy (anonymous):

thanks darling (: